# Multimodal Fusion and Vision-Language Models: A Survey for Robot Vision

### v1

@@ 7 Conclusions
Multimodal fusion and vision-language models have significantly advanced the capabilities of robot vision, enabling enhanced perception, reasoning, and interaction in complex environments. We systematically reviewed their applications across key robotic vision tasks, analyzed the evolution of related technologies, and examined commonly used datasets and evaluation benchmarks. Despite these advancements, challenges such as data heterogeneity, low-quality annotations, and real-time computational constraints persist. Future research should focus on improving cross-modal alignment, developing efficient training and inference strategies, and leveraging self-supervised learning to enhance adaptability. By addressing these challenges, multimodal fusion will continue to drive progress in robotic vision, enabling more autonomous, efficient, and context-aware robotic systems.

### v3

@@ 7 Conclusions
This paper systematically explores the critical role of VLMs in robot vision. Focusing on typical tasks such as semantic understanding, 3D Object detection, embodied navigation, and robotic manipulation, we summarize how different fusion paradigms (e.g., encoder-decoder frameworks, attention mechanisms, graph neural networks, etc.) integrate across modalities such as vision and language, depth, and point clouds. We also conduct a detailed comparison of traditional methods and the evolutionary paths and applicability of recent large-model-based approaches. Additionally, we systematically reviewed relevant datasets and evaluation benchmarks, revealing current imbalances in research across methodological, data, and task dimensions. We have identified the following three key research findings: First, cross-modal alignment strategies are critical to determining the upper limit of a robot’s perception system performance, which is currently constrained by modal differences and mismatches in semantic granularity; Second, lightweight, multi-stage adaptation mechanisms are a practical necessity for deploying large-scale vision-language models on robot platforms, especially in scenarios with limited edge computing resources; Third, task-oriented and online adaptation fusion strategies are increasingly becoming a research focus, aiming to drive vision-language models to dynamically adjust their perception and decision-making paths through scene cognition, thereby supporting robot response capabilities under high-frequency, highly variable target conditions. Although current multimodal and VLM methods perform well on standardized benchmark tasks, they still face numerous challenges on actual robot platforms, including modal incompleteness caused by sensor heterogeneity, semantic bias caused by sparse or low-quality perception inputs, and requirements for latency, robustness, and interpretability during task execution. Future research should focus on introducing structured spatial modeling and memory mechanisms to enhance spatial reasoning, strengthening system interpretability and ethical adaptability, and incorporating brain-inspired modeling principles to develop cognitive VLM architectures with long-term learning capabilities, thereby driving the development of more autonomous, efficient, and explainable intelligent robotic systems.
Acknowledgement
This work was supported by National Natural Science Foundation of China (Nos. 62376271, U21A20515, U22B2034, 62572059, 62172416, 62365014, 62271074 and 62262043), Beijing Natural Science Foundation (JQ23014, L231013) and the Open Project Program of State Key Laboratory of Virtual Reality Technology and Systems, Beihang University (No. VRLAB2025B03).